# An Efficient Iterative Decoupling Method for Thermo-Poroelasticity Based on a Four-Field Formulation

### v2

@@ Acknowledgments
The work of M. Cai is partially supported by the NIH-RCMI award (Grant No. 347U54MD013376), the Army Research Office award (W911NF2310004), and the affiliated project award from the Center for Equitable Artificial Intelligence and Machine Learning Systems (CEAMLS) at Morgan State University (project ID 02232301). The work of J. Li and Z. Li are partially supported by the Shenzhen Sci-Tech Fund No. RCJC20200714114556020, Guangdong Basic and Applied Research Fund No. 2023B1515250005. The work of Q. Liu is partially supported by the NSFC (12271186, 12271178), the Guangdong Basic and Applied Basic Research Foundation (2022B1515120009), and the Science and Technology Program of Shenzhen, China(20231121110406001).

### v3

@@ Acknowledgments
The authors thank two reviewers for their constructive suggestions and comments. The work of M. Cai is partially supported by Army Research Office award W911NF-23-1-0004, the NIH-RCMI award (Grant No. 347U54MD013376), and the affiliated project award from the Center for Equitable Artificial Intelligence and Machine Learning Systems (CEAMLS) at Morgan State University (project ID 02232301). The work of J. Li and Z. Li are partially supported by the Shenzhen Sci-Tech Fund No. RCJC20200714114556020, Guangdong Basic and Applied Research Fund No. 2023B1515250005. The work of Q. Liu is partially supported by the NSFC (12271186, 12271178), the Guangdong Basic and Applied Basic Research Foundation (2022B1515120009), and the Science and Technology Program of Shenzhen, China(20231121110406001).